3. Determine whether a quadratic model exist for the set of values. If so, write the model. Show your work.
f(-2)=16, f(0)=0, f(1)=4

Answer by ewatrrr(24785)

y = ax^2 + bx + c
(0,0)
c = 0
........
y = x^2 + bx
(1,4)
4 = 1 + b
3 = b
....
y = x^2 + 3x
(-2, 16)
16 ≠ 4 -6